# Fact, Fetch, and Reason: A Unified Evaluation of Retrieval-Augmented Generation

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated significant performance improvements across various cognitive tasks. An emerging application is using LLMs to enhance retrieval-augmented generation (RAG) capabilities. These systems require LLMs to understand user queries, retrieve relevant information, and synthesize coherent and accurate responses. Given the increasing real-world deployment of such systems, comprehensive evaluation becomes crucial. To this end, we propose FRAMES (Factuality, Retrieval, And reasoning MEasurement Set), a high-quality evaluation dataset designed to test LLMs’ ability to provide factual responses, assess retrieval capabilities, and evaluate the reasoning required to generate final answers. While previous work has provided datasets and benchmarks to evaluate these abilities in isolation, FRAMES offers a unified framework that provides a clearer picture of LLM performance in end-to-end RAG scenarios. Our dataset comprises challenging multi-hop questions that require the integration of information from multiple sources. We present baseline results demonstrating that even state-of-the-art LLMs struggle with this task, achieving 0.40 accuracy with no retrieval. The accuracy is significantly improved with our proposed multi-step retrieval pipeline, achieving an accuracy of 0.66 (>50% improvement). We hope our work will help bridge evaluation gaps and assist in developing more robust and capable RAG systems.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have significantly enhanced their capabilities across various natural language processing tasks, especially in systems that demand both factual accuracy and sophisticated reasoning for complex queries (Zhao et al., 2023). Retrieval-augmented generation (RAG) techniques (Lewis et al., 2020; Fan et al., 2019; Guu et al., 2020) have become a powerful approach by leveraging the strengths of retrieval systems and the generative capabilities of LLMs. These techniques are particularly effective for tasks requiring multi-hop reasoning, factual grounding, and synthesizing information from diverse knowledge domains (Gao et al., 2023). However, despite this progress, the evaluation of RAG systems remains fragmented and insufficient, as existing benchmarks typically assess components like retrieval, factual correctness, and reasoning in isolation (Yu et al., 2024a). This piecemeal approach fails to capture the holistic performance of these systems in real-world applications (Yu et al., 2024b).
To bridge this gap, we introduce a novel evaluation framework, FRAMES  [^1] Dataset link : https://huggingface.co/datasets/google/frames-benchmark (Factuality, Retrieval, And reasoning MEasurement Set), designed to rigorously test LLMs on all three core capabilities—fact retrieval, reasoning across multiple constraints, and accurate synthesis of information into coherent responses. Unlike existing datasets such as TruthfulQA (Lin et al., 2021), HotpotQA (Yang et al., 2018b), or GSM8k (Cobbe et al., 2021), which focus on isolated aspects of LLM performance, FRAMES provides an integrated evaluation that challenges models across these dimensions simultaneously. This approach offers a more accurate reflection of how these systems perform as end-to-end reasoning solutions, especially in scenarios requiring multi-document retrieval and complex reasoning. For instance, a sample from our dataset is "How many years earlier would Punxsutawney Phil have to be canonically alive to have made a Groundhog Day prediction in the same state as the US capitol?" This requires the model to perform temporal and numerical reasoning after retrieving the critical articles needed to answer the question.
@@
  • We introduce FRAMES, a novel dataset of 824 test samples designed to evaluate LLMs’ ability to retrieve and reason across multiple documents in a unified framework.
  • We propose a multi-step retrieval and reasoning framework that compels models to iteratively retrieve and reason, significantly enhancing their performance on complex queries.
Dataset Factuality Retrieval Reasoning Multi-Hop/Step Temporal Disambiguation
FRAMES
@@
Multihop-RAG(Tang & Yang, 2024)
@@ 2 FRAMES
FRAMES (Factuality, Retrieval, And reasoning MEasurement Set) is an evaluation set of 824 questions designed to provide an end-to-end evaluation of Retrieval Augmented Generation (RAG) systems. It assesses three key components of a RAG system: Factuality, Retrieval, and Reasoning. Unlike most existing datasets and benchmarks that evaluate each of these RAG components in isolation, FRAMES offers a comprehensive test bed to gain a clear understanding of the overall quality of RAG systems(Lin et al., 2022; Yang et al., 2018a; Welbl et al., 2017). This holistic approach allows for a more accurate reflection of how these systems perform in real-world scenarios. In this section, we first detail our data collection process, which involved both synthetic data generation attempts and human annotation. Next, we present the dataset statistics, showcasing the diversity of topics and reasoning types covered. Finally, we outline the rigorous quality checks implemented to ensure the dataset’s reliability and challenging nature. By providing this end-to-end evaluation framework, FRAMES aims to bridge the gap in existing benchmarks and foster the development of more robust and efficient RAG systems.
Task Instruction Prompt for Human Annotation Task Description Create n factoid questions that demand multi-hop reasoning based on information found across multiple Wikipedia articles. These questions should ideally have a single, unambiguous answer and may optionally incorporate elements of challenging reasoning. Here’s a breakdown of the key terms and their relationships: Factoid Questions: These are trivia-style questions with a single, clearly defined, and factually correct answer. Multi-hop Reasoning: This refers to the core requirement that answering the questions necessitates combining information from different sections within multiple Wikipedia articles. The example provided ("What is the name of the river that flows through the city where the Eiffel Tower is located?") highlights how this differs from a simple factoid question ("What is the capital of France?"). Challenging Reasoning: This aspect encourages the inclusion of questions that go beyond simple information retrieval and demand critical thinking. This can be achieved through various question types like: Numerical Reasoning: Involving counting, comparisons, or calculations. <examples> Tabular Reasoning: Involving statistics found in tables / info boxes in wikipedia. <examples> Multiple Constraints: Questions involving multiple constraints, whose intersection points towards a unique answer. <examples> Temporal Reasoning : Questions involving reasoning through timelines. <examples> Post processing: This involves requiring the answerer to perform specific post-processing steps after all necessary facts have been gathered. <examples> Additional Requirements: Wikipedia Articles: The information used to answer the questions must be sourced from Wikipedia articles. Standalone and Context-Independent: Questions should be understandable without requiring additional information or context. Single, Unambiguous Answer: Each question should have only one correct answer, leaving no room for ambiguity. Avoid boolean questions (yes/no questions) that can be answered with a simple "yes" or "no." <examples>
Figure 1: Task instruction provided to human annotators to generate samples for FRAMES.
@@
We start our data collection process with synthetic dataset generation to explore a potentially cost-effective alternative to expensive human annotation. We prompt a state-of-the-art LLM with instructions to use multiple articles to generate questions that would require information from multiple articles to answer. The prompt (shown in Figure 6 in Appendix A) takes as input the number of articles provided to generate questions. However, we observed significant issues with this approach. While the LLMs were able to generate coherent questions, there was a high proportion of hallucinated questions and answers (>30%). Additionally, the LLM struggled to generate questions that strictly required more than four articles. To evaluate the potential of this approach, we manually cleaned the hallucinated questions and answers from the obtained set. We then evaluated the same LLM on these cleaned questions and obtained an accuracy of $\sim$32%, suggesting that the legitimate questions generated by LLMs were indeed challenging for state-of-the-art models. There are two key takeaways from our experimentation with synthetic data generation: (1) Synthetic test data requires heavy manual cleaning before usage, which suggests that we will need to rely on human annotations instead of LLMs to generate the final evaluation set; and (2) models performed significantly poorly on the correct test samples we tested on, suggesting that the instruction to create questions can be used to generate a challenging evaluation set.
Human annotation.
@@
  • Addressing contamination issues: To mitigate concerns about potential contamination due to Wikipedia articles being in LLM training sets, we designed questions that require additional reasoning and operations beyond simple fact retrieval. For instance, the question "How many years earlier would Punxsutawney Phil have to be canonically alive to have made a Groundhog Day prediction in the same state as the US capitol?" requires not only fact extraction but also additional calculations.
@@ 3.1 Single-Step Evaluations
In this set of experiments, we evaluate the model using several baseline prompting methods on our test set to understand how well existing LLMs perform. Specifically, we experiment with three baseline approaches: (1) Naive Prompt: This is a straightforward approach where we simply ask the question to the model and evaluate if the model’s response without search retrieval contains the correct answer. (2) BM25-Retrieved Prompt (n_docs): This approach augments the question with the top n_docs documents having the highest BM25 score (Robertson et al., 1995) retrieved from a Wikipedia data dump [^2] https://www.tensorflow.org/datasets/catalog/wikipedia#wikipedia20230601en_default_config. The BM25 score is computed between the question and every article in the Wikipedia dump, after which the top n_docs with the highest scores are added to the prompt. The motivation behind this approach is to observe improvements in model performance when relevant articles are added to the context. This is denoted as BM25-R (n_doc) in the results. (3) Oracle Prompt: This prompt includes the question along with all the ground truth Wikipedia articles used by the human annotators to generate the question. The performance of the Oracle Prompt provides the upper bound of model performance in the case of a perfect retrieval system that is able to extract all the relevant articles.
Experiment Setup.
For the experiments, we use Gemini-Pro-1.5-0514 (Google, 2024b), Gemini-Flash-1.5-0514 (Google, 2024a), Gemma2-27b (Gemma et al., 2024), and Gemma2-9b (Gemma et al., 2024) as the state-of-the-art LLM since they have shown great performance on several public benchmarks. Since the gold answers to questions in the dataset are free-form tokens instead of choices from multiple-choice answers, we use an LLM to evaluate if the outcome from the LLM under evaluation matches the gold answer, using the prompt shown in Figure 7 in Appendix B. This auto-rating mechanism was tested against human evaluations, in which the LLM-based evaluation showed strong alignment with human annotations (accuracy: 0.96 and Cohen’s Kappa: 0.889 for Gemini-Pro-1.5-0514 as autorating LLM), making LLM-based evaluation a suitable approach to evaluate the correctness of model responses.
Figure 3: Accuracy of Gemini-Pro-1.5-0514 across different reasoning types in our test set. The results indicate superior performance on logical and temporal reasoning tasks, with notable deficiencies in numerical, tabular, and post-processing reasoning. The substantial performance improvements observed with oracle information underscore the critical role of relevant contextual information in enhancing model accuracy across all reasoning categories.
@@
Based on results shown in Table 3, we observe that naive prompting attains a performance of $\sim$40with gradual increases when including BM25 retrieved articles for Gemini-Pro-1.5-0514. The model achieves an accuracy of $\sim$45when the number of documents in the context is 2, and $\sim$47when double the number of articles are added to the context. These improvements demonstrate the room for enhancement when the model is able to retrieve relevant articles required to answer the question. The core reason behind these improvements is the improvement in recall in the articles present in context which increased from 0.12 (BM25-R(n_docs = 2) to 0.15 (BM25-R (n_docs = 4)). In addition to these approaches, we observe an accuracy of $\sim$72for Gemini-Pro-1.5-0514 when all the gold Wikipedia articles are provided in the context, which we call Oracle Prompt. Out of $\sim$28samples where the model made errors, $\sim$80of those misclassifications belong to numerical, tabular, and post-processing categories. Hence, these misclassifications show the reasoning gaps in model performance where even after providing all the relevant facts, the model failed to reason through the different facts to provide a correct answer to the question. The accuracies obtained by the Naive Prompt and Oracle Prompt can be considered as the lower bound (when no relevant articles were provided to the model) and upper bound (when all relevant articles were provided to the model) of model performances on FRAMES. This pattern can also be seen in Figure 3 where we plotted accuracy for each reasoning type and observe that the model performed the lowest in numerical, post-processing, and tabular reasoning tasks. We also observe that adding BM25 retrieved articles primarily helped with questions requiring reasoning through multiple constraints ($\sim$8improvement) and post-processing ($\sim$10improvement). This aligns well with the fact that providing more relevant articles helps in obtaining facts for each constraint, leading to improvements in performance. We take these learnings and experiment with a more complicated setup where the model is asked to find answer to questions through multiple iterations instead of a single step.
3.2 Multi-Step Evaluations
Baselines G-Pro-1.5 G-Flash-1.5 Gemma2-27b Gemma2-9b
Naive Prompt 0.408 0.263 0.308 0.051
BM25-R (n__\__doc=2) 0.452 0.288 - -
BM25-R (n__\__doc=4) 0.474 0.315 - -
Oracle Prompt 0.729 0.665 - -
Table 3: This table presents the accuracy performance of Gemini-Pro-1.5-0514 (G-Pro-1.5), Gemini-Flash-1.5-0514 (G-Flash-1.5), Gemma2-27b, and Gemma2-9b on our proposed evaluation dataset. Please note that the performance of Gemma models is not reported for cases requiring longer context due to the small maximum context length of the model.
@@ 3.2 Multi-Step Evaluations
Based on the findings from the previous experiment with single-step evaluations, where we observed an increase in performance when related articles are added to the context, wew were led to explore a setting where the model is compelled to plan its search for relevant Wikipedia articles in order to find answers. More specifically, we design a pipeline where the model is asked a question along with the instruction to generate $k$ search queries which are then used to extract the top-n_docs Wikipedia articles with the highest BM25 scores. These documents are then appended to the context. This process of query generation and retrieved article augmentation is carried forward for $n$ steps. Once the $n$ steps of retrieval are completed, the model is asked to answer the question based on the articles appended in the context, as shown in Algorithm 1. We conduct two sets of experiments here: (1) Vanilla with no explicit planning instructions, and (2) With search planning instructions to help the model navigate the search process efficiently. To implement this pipeline, we used the simplest document retrieval component which is essentially an index of Wikipedia pages, where the articles with the highest BM25 scores for each query are returned to the LLM and added to the context. This retrieval component is used instead of making direct calls to an online search engine for two reasons: (1) We would like to keep the retrieval system constant to clearly evaluate the search planning capability of the LLMs instead of the retrieval system’s capability in returning the most relevant articles, and (2) The BM25-based retrieval system makes our pipeline reproducible and limits the search space to Wikipedia pages only, as the questions were generated from Wikipedia articles only.

### v2

@@ Abstract
Large Language Models (LLMs) have shown significant improvements across cognitive tasks, with an emerging application in enhancing retrieval-augmented generation (RAG) capabilities. These systems require LLMs to understand queries, retrieve relevant information, and synthesize accurate responses. Given their increasing real-world deployment, comprehensive evaluation is crucial. We propose FRAMES (Factuality, Retrieval, And reasoning MEasurement Set), a high-quality dataset designed to test LLMs’ factual responses, retrieval capabilities, and reasoning in generating final answers. Unlike previous work evaluating these abilities in isolation, FRAMES offers a unified framework for assessing LLM performance in end-to-end RAG scenarios. Our dataset comprises challenging multi-hop questions requiring integration of information from multiple sources. Baseline results show that even state-of-the-art LLMs struggle, achieving 0.408 accuracy without retrieval. However, our proposed multi-step retrieval pipeline significantly improves accuracy to 0.66 (>50% improvement). We aim to bridge evaluation gaps and assist in developing more robust RAG systems.
Fact, Fetch, and Reason: A Unified Evaluation of
Retrieval-Augmented Generation
Satyapriya Krishnathanks: Work done as an intern at Google.1, Kalpesh Krishnathanks: Equal contribution2, Anhad Mohananey2, Steven Schwarcz2, Adam Stambler2, Shyam Upadhyay2, Manaal Faruqui2 1Harvard University, 2Google, Inc. Correspondence: skrishna@g.harvard.edu
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have significantly enhanced their capabilities across various natural language processing tasks, especially in systems that demand both factual accuracy and sophisticated reasoning for complex queries (Zhao et al., 2023; Krishna et al., 2024). Retrieval-augmented generation (RAG) techniques (Lewis et al., 2020; Fan et al., 2019; Guu et al., 2020) have become a powerful approach by leveraging the strengths of retrieval systems and the generative capabilities of LLMs. These techniques are particularly effective for tasks requiring multi-hop reasoning, factual grounding, and synthesizing information from diverse knowledge domains (Gao et al., 2023). However, despite this progress, the evaluation of RAG systems remains fragmented and insufficient, as existing benchmarks typically assess components like retrieval, factual correctness, and reasoning in isolation (Yu et al., 2024a). This piecemeal approach fails to capture the holistic performance of these systems in real-world applications (Yu et al., 2024b).
Figure 1: An example from the FRAMES dataset, highlighting the core capabilities needed by a system (Factuality, Retrieval, Reasoning) to answer the question.
To bridge this gap, we introduce a novel evaluation framework, FRAMES  [^1] Dataset link : https://huggingface.co/datasets/google/frames-benchmark (Factuality, Retrieval, And reasoning MEasurement Set), designed to rigorously test LLMs on all three core capabilities—fact retrieval, reasoning across multiple constraints, and accurate synthesis of information into coherent responses. Unlike existing datasets such as TruthfulQA (Lin et al., 2021), HotpotQA (Yang et al., 2018a), or GSM8k (Cobbe et al., 2021), which focus on isolated aspects of LLM performance, FRAMES provides an integrated evaluation that challenges models across these dimensions simultaneously. This approach offers a more accurate reflection of how these systems perform as end-to-end reasoning solutions, especially in scenarios requiring multi-document retrieval and complex reasoning. For example, our dataset includes questions like: "How many years earlier would Punxsutawney Phil have to be canonically alive to have made a Groundhog Day prediction in the same state as the US capitol?" This demands temporal and numerical reasoning based on information from multiple retrieved articles. Figure 1 provides another such example.
@@
Dataset Factuality Retrieval Reasoning Multi-Hop or Multi-Step Temporal Disambiguation
FRAMES (our work)
@@
Multihop-RAG(Tang and Yang, 2024)
@@
  • We introduce FRAMES, a novel dataset of 824 test samples designed to evaluate LLMs’ ability to retrieve and reason across multiple documents in a unified framework. We plan to open-source this dataset for public use.
  • We propose a multi-step retrieval and reasoning framework that compels models to iteratively retrieve and reason, significantly enhancing their performance on complex queries, from an accuracy of 0.408 with single-step inference to 0.66 with multi-step retrievals.
@@ 2 FRAMES
FRAMES (Factuality, Retrieval, And reasoning MEasurement Set) is an evaluation set of 824 questions designed to provide an end-to-end evaluation of Retrieval Augmented Generation (RAG) systems. It assesses three key components of a RAG system: Factuality, Retrieval, and Reasoning. Unlike most existing datasets and benchmarks that evaluate each of these RAG components in isolation, FRAMES offers a comprehensive test bed to gain a clear understanding of the overall quality of RAG systems (Lin et al., 2022; Yang et al., 2018b; Welbl et al., 2017). This holistic approach allows for a more accurate reflection of how these systems perform in real-world scenarios. In this section, we first detail our data collection process, which involved both synthetic data generation attempts and human annotation. Next, we present the dataset statistics, showcasing the diversity of topics and reasoning types covered. Finally, we outline the rigorous quality checks implemented to ensure the dataset’s reliability and challenging nature. By providing this end-to-end evaluation framework, FRAMES aims to bridge the gap in existing benchmarks and foster the development of more robust and efficient RAG systems.
@@
We start our data collection process with synthetic dataset generation to explore a potentially cost-effective alternative to expensive human annotation. We prompt a state-of-the-art LLM with instructions to use multiple articles to generate questions that would require information from multiple articles to answer. The prompt (shown in Figure 5 in the Appendix) takes as input the number of articles provided to generate questions. However, we observed significant issues with this approach. While the LLMs were able to generate coherent questions, there was a high proportion of hallucinated questions and answers (>30%). Additionally, the LLM struggled to generate questions that strictly required more than four articles. To evaluate the potential of this approach, we manually cleaned the hallucinated questions and answers from the obtained set. We then evaluated the same LLM on these cleaned questions and obtained an accuracy of $\sim$32%, suggesting that the legitimate questions generated by LLMs were indeed challenging for state-of-the-art models. There are two key takeaways from our experimentation with synthetic data generation: (1) Synthetic test data requires heavy manual cleaning before usage, which suggests that we will need to rely on human annotations instead of LLMs to generate the final evaluation set; and (2) models performed significantly poorly on the correct test samples we tested on, suggesting that the instruction to create questions can be used to generate a challenging evaluation set.
@@
  • Addressing data contamination issues: To mitigate concerns about potential contamination due to Wikipedia articles being in LLM training sets, we designed questions that require additional reasoning and operations beyond simple fact retrieval. For instance, the question "How many years earlier would Punxsutawney Phil have to be canonically alive to have made a Groundhog Day prediction in the same state as the US capitol?" requires not only fact extraction but also additional calculations.
@@ 3 Empirical Analysis
Baselines G-Pro-1.5 G-Flash-1.5 Gemma2-27b LLama3.2-3B-I Qwen2.5-3B-I
Date of Release 5/14/24 5/14/24 6/27/24 9/25/24 9/19/24
Naive Prompt 0.408 0.263 0.308 0.115 0.095
BM25-R (n__\__doc=2) 0.452 0.288
BM25-R (n__\__doc=4) 0.474 0.315
Oracle Prompt 0.729 0.665
Table 3: This table presents the accuracy performance of Gemini-Pro-1.5-0514 (G-Pro-1.5), Gemini-Flash-1.5-0514 (G-Flash-1.5), Gemma2-27b, LLama3.2-3B-I and Qwen2.5-3B-I on our proposed evaluation dataset. Please note that the performance of Gemma, Llama, and Qwen models is not reported for cases requiring longer context due to the small maximum context length of the model.
@@ 3.1 Single-Step Evaluations
In this set of experiments, we evaluate the model using several baseline prompting methods on our test set to understand how well existing LLMs perform. Specifically, we experiment with three baseline approaches: (1) Naive Prompt: This is a straightforward approach where we simply ask the question to the model and evaluate if the model’s response without search retrieval contains the correct answer. (2) BM25-Retrieved Prompt (n_docs): This approach augments the question with the top n_docs documents having the highest BM25 score (Robertson et al., 1995) retrieved from a Wikipedia data dump [^2] https://tinyurl.com/36jxum2y. The BM25 score is computed between the question and every article in the Wikipedia dump, after which the top n_docs with the highest scores are added to the prompt. The motivation behind this approach is to observe improvements in model performance when relevant articles are added to the context. This is denoted as BM25-R (n_doc) in the results. (3) Oracle Prompt: This prompt includes the question along with all the ground truth Wikipedia articles used by the human annotators to generate the question. The performance of the Oracle Prompt provides the upper bound of model performance in the case of a perfect retrieval system that is able to extract all the relevant articles.
Experimental and Autorater setup.
For the experiments, we use Gemini-Pro-1.5-0514 (Google, 2024b), Gemini-Flash-1.5-0514 (Google, 2024a), Gemma2-27b (Gemma et al., 2024), LLama3.2-3B-I (Meta, 2024), and Qwen2.5-3B-I (Qwen, 2024) as the state-of-the-art LLM since they have shown great performance on several public benchmarks. Since the gold answers to questions in the dataset are free-form tokens instead of choices from multiple-choice answers, we use an LLM to evaluate if the outcome from the LLM under evaluation matches the gold answer, using the prompt shown in Figure 6 in Appendix. This auto-rating mechanism was tested against human evaluations, in which the LLM-based evaluation showed strong alignment with human annotations (accuracy: 0.96 and Cohen’s Kappa: 0.889 for Gemini-Pro-1.5-0514 as autorating LLM), making LLM-based evaluation a suitable approach to evaluate the correctness of model responses.
(a) Single-Step Evaluations
(b) Multi-Step Evaluations
Figure 3: This plot shows model performance for each reasoning type in single-step (a) and multi-step (b) evaluations. Reasoning types: MC (Multiple Constraints), NR (Numerical Reasoning), PP (Post Processing), TR (Tabular Reasoning), and TeR (Temporal Reasoning). (a) Gemini-Pro-1.5-0514’s accuracy across reasoning types in single-step evaluations. Results show better performance in logical and temporal reasoning, with weaknesses in numerical, tabular, and post-processing reasoning. Oracle information significantly improves accuracy across all categories. (b) Gemini-Pro-1.5-0514’s accuracy in multi-step evaluations. Multi-step search planning increases performance for all reasoning types, with numerical reasoning surpassing oracle performance.
@@
Based on results shown in Table 3, we observe that naive prompting attains a performance of 0.408 with gradual increases when including BM25 retrieved articles for Gemini-Pro-1.5-0514. The model achieves an accuracy of 0.452 when the number of documents in the context is 2, and 0.474 when double the number of articles are added to the context. These improvements demonstrate the room for enhancement when the model is able to retrieve relevant articles required to answer the question. The core reason behind these improvements is the improvement in recall in the articles present in context which increased from 0.12 (BM25-R(n_docs = 2) to 0.15 (BM25-R (n_docs = 4)). In addition to these approaches, we observe an accuracy of 0.729 for Gemini-Pro-1.5-0514 when all the gold Wikipedia articles are provided in the context, which we call Oracle Prompt. Out of 27% samples where the model made errors, $\sim$80of those misclassifications belong to numerical, tabular, and post-processing categories. Hence, these misclassifications show the reasoning gaps in model performance where even after providing all the relevant facts, the model failed to reason through the different facts to provide a correct answer to the question. The accuracies obtained by the Naive Prompt and Oracle Prompt can be considered as the lower bound (when no relevant articles were provided to the model) and upper bound (when all relevant articles were provided to the model) of model performances on FRAMES. This pattern can also be seen in Figure 3(a) where we plotted accuracy for each reasoning type and observe that the model performed the lowest in numerical, post-processing, and tabular reasoning tasks. We also observe that adding BM25 retrieved articles primarily helped with questions requiring reasoning through multiple constraints ($\sim$8improvement) and post-processing ($\sim$10improvement). This aligns well with the fact that providing more relevant articles helps in obtaining facts for each constraint, leading to improvements in performance. We take these learnings and experiment with a more complicated setup where the model is asked to find answer to questions through multiple iterations instead of a single step.
@@
3.2 Multi-Step Evaluations
Based on the findings from the previous experiment with single-step evaluations, where we observed an increase in performance when related articles are added to the context, we were led to explore a setting where the model is compelled to plan its search for relevant Wikipedia articles in order to find answers. More specifically, we design a pipeline where the model is asked a question along with the instruction to generate $k$ search queries which are then used to extract the top-n_docs Wikipedia articles with the highest BM25 scores. These documents are then appended to the context. This process of query generation and retrieved article augmentation is carried forward for $n$ steps. Once the $n$ steps of retrieval are completed, the model is asked to answer the question based on the articles appended in the context, as shown in Algorithm 1. We conduct two sets of experiments here: (1) Vanilla with no explicit planning instructions, and (2) With search planning instructions to help the model navigate the search process efficiently. To implement this pipeline, we used the simplest document retrieval component which is essentially an index of Wikipedia pages, where the articles with the highest BM25 scores for each query are returned to the LLM and added to the context. This retrieval component is used instead of making direct calls to an online search engine for two reasons: (1) We would like to keep the retrieval system constant to clearly evaluate the search planning capability of the LLMs instead of the retrieval system’s capability in returning the most relevant articles, and (2) The BM25-based retrieval system makes our pipeline reproducible and limits the search space to Wikipedia pages only, as the questions were generated from Wikipedia articles only.